OpenStudy (anonymous):

Under what circumstance would the sum of two fractions be easier to compute than the product of the same two fractions?

OpenStudy (anonymous):

i guess when they have the same denominator

OpenStudy (turingtest):

\[\frac ac+\frac bc=\frac{a+b}c\]

OpenStudy (anonymous):

thanks, sound good to me.

OpenStudy (phi):

but for the multiplication to be difficult you want a,b and c to be large numbers so for example \[ \frac{-12345}{78901234} + \frac{12345}{78901234} \] is much easier than \[ \frac{-12345}{78901234} \cdot \frac{12345}{78901234} \]

OpenStudy (phi):

as compared to \[ \frac{1}{2}+\frac{1}{2} \] and \[ \frac{1}{2} \cdot \frac{1}{2} \] Both seem easy.